### v1

@@ 3 Worst Case Bounds on Certificate Size
Let $\{(x_{1},y_{1}),\dots,(x_{s{b}},y_{s{b}})\}$ be the largest $b$-robust hollow star for $\mathcal{H}$, with corresponding weights $w_{1},\dots,w_{s{b}}$. Let $i^{*}$ be such that $w_{i^{*}}=b+1$. We will set the training dataset $S=\{(x_{i},y_{i}):i\neq i^{*}\}$, test point to be $x_{i^{\star}}$ and test label to be $1-y_{i^{\star}}$, and argue that $S$ is a minimal $b$-robust certificate for $(x_{i^{\star}},1-y_{i^{\star}})$.
Let $T$ be the weighted sequence $\{(x_{1},y_{1},w_{1}),\dots,(x_{s{b}},y_{s{b}},w_{s{b}})\}$. By definition of the $b$-robust hollow star, $T^{\prime}=T\setminus\{(x_{i^{*}},y_{i^{*}},w_{i^{*}})\}$ is $b$-robustly realizable by $\mathcal{H}$. But note that all the weights in $T^{\prime}$ are equal to $1$. This means that there exists $h\in\mathcal{H}$ such that $\sum_{i\in[s{b}],i\neq i^{*}}\mathds{1}[h(x_{i})\neq y_{i}]\leq b$. Namely, $S$ is $b$-robustly realizable by $\mathcal{H}$.
Next, consider any $h\in\mathcal{H}$ that satisfies $\sum_{i\in[s{b}],i\neq i^{*}}\mathds{1}[h(x_{i})\neq y_{i}]\leq b$. Then, it must be the case that $h(x_{i^{*}})=1-y_{i^{*}}$; otherwise, the weighted sequence
$T=\{(x_{1},y_{1},w_{1}),\dots,(x_{i^{*}},y_{i^{*}},w_{i^{*}}),\dots,(x_{s{b}},y_{s{b}},w_{s{b}})\}$
would be $b$-robustly realizable by $\mathcal{H}$, which would contradict that $\{(x_{1},y_{1}),\dots,(x_{s{b}},y_{s{b}})\}$ is a $b$-robust hollow star. Thus, $S$ is a $b$-robust certificate for $(x_{i^{*}},1-y_{i^{*}})$.
@@
Let us revisit the examples of halfspaces and singletons considered earlier for the setting with no corruptions. For $d$-dimensional halfspaces passing through the origin (Example 3.1), our certificate based on Carathéodory’s theorem had size $d$. Indeed, this class has hollow star number equal to $d+1$, which means that this certificate size is optimal in general. On the other hand, the class of singletons on a domain of size $n$ has hollow star number equal to $n$. This validates why we were unable to obtain a certificate of size smaller than $n-1$ in Example 3.2.
Thus, we have shown that in general, the $b$-robust hollow star number optimally characterizes minimum reliable certificate size. While Claim 3.6 relates the $b$-robust hollow star to the hollow star number (which is known for many natural classes, e.g., see the examples in Section 2.1 in [5]), it merely gives a lower bound. It would be interesting to see if an upper bound can also be obtained, and more generally to chart out the $b$-robust hollow star for natural classes.
For example, we can show that the $b$-robust hollow star number for singletons on a domain of size $n$ is exactly $(b+1)(n-1)+1$ (Section A.2). In addition, for $d$-dimensional halfspaces, one can use the same reasoning as in Example 3.1, together with the Tolerance Carathéodory theorem (Thereom 4.1 in [25], Theorem 2 in [31]) to obtain certificates of size $<(d+b)^{O(b)}$ for test points in the $b$-robust agreement of the training data. From Theorem 3.7, we then know that the $b$-robust hollow star number of halfspaces is at most $(d+b)^{O(b)}$ (the lower bound from Claim 3.6 is only $\Omega(bd)$.) One way to close the gap for halfspaces would be to show a lower bound for the Tolerance Carathéodory theorem. The simplest phrasing of this for $b=1$ amounts to the following purely convex-analytic question:
@@
Such a result exists for the Tolerance Helly Theorem [25, Theorem 3.2].
@@ 4 Worst-case Distributional Bounds on Certificate Size
A first observation is that if the distribution $\mathcal{D}$ is discrete, and the test point $x$ belongs to the support of $\mathcal{D}$, we will eventually observe more than $12b+$ 1 copies of it in $S$. With a corruption budget of $b$, an adversary can potentially corrupt the labels on $b$ of these copies; however, $b+1$ copies with the true label $h^{\star}(x)$ still remain in $S$. We can use these copies, all of which have the true label, to certify that $x$ should be labeled as $h^{\star}(x)$, since any hypothesis that labels $x$ differently makes strictly more than $b$ mistakes on these copies. Furthermore, this certificate size is optimal—any certificate of smaller size is $b$-robustly realizable by every hypothesis in the class.
@@
Let $m^{\prime}$ be such that $\operatorname*{\mathbb{P}}_{S\sim\mathcal{D}^{m^{\prime}}}[(x,h^{\star}(x))\text{ in agreement region of }S]=\delta^{\prime}$ for some $\delta^{\prime}>0$; such an $m^{\prime}$ exists by assumption that $(x,h^{\star}(x))$ belongs to the agreement region of a drawn sample $S$ eventually. Consider drawing a sample $S$ of size $(2b+1)nm^{\prime}/\delta^{\prime}$, and let $S_{1},S_{2},\dots,S_{(2b+1)n/\delta^{\prime}}$ denote independent partitions of $S$ into chunks of size $m^{\prime}$. The expected number of chunks that satisfy that $(x,h^{\star}(x))$ is in the agreement region of the chunk is $(2b+1)n$. By a Chernoff bound, the probability that at least $12b+$ 1 chunks have $(x,h^{\star}(x))$ in their agreement region is at least $1-e^{-\Omega(bn)}$. Now, with a corruption budget of size $b$, an adversary can insert a corruption in at most $b$ chunks. Even so, the remaining $\geq b+1$ chunks continue to have $(x,h^{\star}(x))$ in their agreement region. Without loss of generality, suppose $S_{1},\dots,S_{b+1}$ are all uncorrupted. Then, from Theorem 3.7 (with $b=0$), we know that each of $S_{1},\dots,S_{b+1}$ will contain a certificate for $(x,h^{\star}(x))$ of size at most $s_{0}-1$; let these certificates be $C_{1},\dots,C_{b+1}$. The final $b$-robust certificate is simply a concatenation of $C_{1},\dots,C_{b+1}$.
@@
and this $h$ must label $x_{i}$ as $-y_{i}$ (by non-realizability of the hollow star). Then, because the certificate only consists of at most $b$ copies of $(x_{i},y_{i})$, it is $b$-robustly realizable by $h$. This contradicts it being a $b$-robustly realizable certificate for $(x_{s_{0}},-y_{s_{0}})$, since $h$ labels $x_{s_{0}}$ as $y_{s_{0}}$.
@@ 5.1 Upper Bound on the Sample Complexity
We observe that the proof of Theorem 5.2 can be adapted to ensure that $S$ robustly certifies multiple test points simultaneously. Concretely, for a class $\mathcal{H}$ having VC dimension $d$, target hypothesis $h^{\star}\in\mathcal{H}$, and marginal distribution $\mathcal{D}$, let $\mathcal{X}_{\varepsilon}=\{x\in\mathcal{X}:\varepsilon_{x}(\mathcal{D},h^{\star})\geq\varepsilon\}$. Observe that any $h$ that mislabels at least one point—say $x_{1}$—in $\mathcal{X}_{\varepsilon}$ satisfies that $h\in\mathcal{H}_{x_{1}}\implies\operatorname*{\mathbb{P}}_{z\sim\mathcal{D}}[h(z)\neq h^{\star}(z)]\geq\varepsilon$ by definition of $\varepsilon_{x_{1}}$. From here, the analysis in the proof of Theorem 5.2 gives us that for a sample $S$ of size $O\left(\frac{b+d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right)$, with probability $1-\delta$, every point in $\mathcal{X}_{\varepsilon}$ belongs to the $b$-robust agreement region of $S$ [^3] One can also analyze the probability mass of $\mathcal{X}_{\epsilon}$, as it is directly related to the disagreement coefficient $\theta$ introduced by [21]. Specifically, it holds that $\operatorname*{\mathbb{P}}[x\in\mathcal{X}_{\varepsilon}]=1-\operatorname*{\mathbb{P}}[x\in\mathcal{X}\setminus\mathcal{X}_{\varepsilon}]\geq 1-\theta\varepsilon$.
@@ 5.2 Tightness of the Upper Bound
To show the tightness of the other two terms, we will work with the following class $\mathcal{H}$. For any given $d\geq 1$, the domain $\mathcal{X}$ of the class will be $\{x_{0},\dots,x_{k}\}$, where we can pick $k$ to be any integer that satisfies $1+\sqrt{\frac{100}{d}}<\frac{1}{2}\log\left(\frac{k}{d}\right)$. [^4] This is for the purpose of instantiating a coupon collector bound, e.g., as in Remark 20, [5]. For example, $k=(100d)^{20}$ is a valid choice. The hypothesis class $\mathcal{H}$ comprises of $\binom{k}{d}+1$ hypotheses. For every $S\subset\{x_{1},\dots,x_{k}\},|S|=d$, there is a hypothesis $h_{S}\in\mathcal{H}$ that labels exactly the set $S$ as 1, and the rest of the domain as $-1$. Additionally, there is a special hypothesis $h^{\star}\in\mathcal{H}$ which satisfies that $h^{\star}(x_{0})=1,h^{\star}(x_{i})=-1,\forall i\in[k]$. That is, $h^{\star}$ only labels $x_{0}$ as 1. Moreover, $h^{\star}$ is the only hypothesis in $\mathcal{H}$ that labels $x_{0}$ as $1$; all the other $\binom{k}{d}$ hypotheses label $x_{0}$ as $-1$. It can be readily verified that the VC dimension of $\mathcal{H}$ is $d$. In what follows, we will set $b=0$ for convenience.
We will set the target hypothesis to be $h^{\star}$, the data distribution $\mathcal{D}$ to be uniform on $\{x_{1},\dots,x_{k}\}$, and the test point $x$ to be $x_{0}$. Crucially, $\mathcal{D}$ has zero mass on $x_{0}$.  [^5] We can also allow a mass of $o(\varepsilon_{x}/(d\log(1/\varepsilon_{x})))$ on $x_{0}$—this ensures $x_{0}$ is not seen in the sample with good chance. Recall that $h^{\star}(x_{0})=1$ and $h^{\star}(x_{i})=-1,\forall i\in[k]$. Since every other hypothesis in $\mathcal{H}$ labels $x_{0}$ as $-1$, and also labels some $d$ points in $\{x_{1},\dots,x_{k}\}$ as 1, we have that $\varepsilon_{x}=\frac{d}{k}$. Now suppose that we obtain a sample of size $m=\frac{d}{2\varepsilon_{x}}\log\left(\frac{1}{\varepsilon_{x}}\right)=\frac{k}{2}\log\left(\frac{k}{d}\right)$. Then, by a coupon collector argument (see Lemma 19 in [5]), we have that with probability at least $\frac{1}{100}>\delta$, the number of distinct elements from $\{x_{1},\dots,x_{k}\}$ that we see in the sample are at most $k-d$. But this means that there is some subset $S\subset\{x_{1},\dots,x_{k}\}$ of $d$ elements whose labels we have not seen. Thus, $h_{S}$ is consistent with the labeled data seen so far, rendering $(x,h^{\star}(x))$ to not be in the agreement region of the sample.
@@ 6 Shorter Certificates by Reweighting
Consider the case when $\mathcal{H}$ is the class of halfspaces in $\mathbb{R}^{d}$ (not necessarily passing through the origin), and $\mathcal{D}$ is uniform on the unit ball $\{x:\|x\|_{2}\leq 1\}$. Suppose that the target halfspace $h^{\star}$ labels every point in the ball positively (i.e., it does not cut through the ball), and that the test point $x$ is at (1/2,0,0,,0(1/2,0,0$\dots$,0( 1 / 2 , 0 , 0 , … , 0). We can verify that $\varepsilon_{x}\leq 2^{-\Omega(d)}$ (as realized by the halfspace $x_{1}<1/2$; see Theorem 2.7 in [4] for the volume of the ball excluded by this halfspace), and so, the upper bound in Theorem 5.2 would suggest that we draw $b\cdot 2^{\Omega(d)}$ samples to ensure that $x$ lies in the $b$-robust agreement region of the samples with constant probability. We might extract a shorter certificate from this sample using the worst-case bound from Theorem 3.7, but even this scales as the $b$-robust hollow star number of $\mathcal{H}$, for which the best upper bound is $(d+b)^{O(b)}$. Consider the following alternate recipe. Suppose we discard any samples from $\mathcal{D}$ that are not contained in the ball of radius $1/2$ centered at $x$. This induces the uniform distribution $\mathcal{D}_{w}$ on the smaller ball, and in expectation, we obtain one sample from $\mathcal{D}_{w}$ for every $2^{d}$ samples from $\mathcal{D}$. Crucially, notice that $\varepsilon_{x}(\mathcal{D}_{w},h^{\star})=1/2$. By Theorem 5.2, $O(b+d)$ samples from $\mathcal{D}_{w}$ are sufficient to ensure that $x$ lies in the $b$-robust agreement region of the samples with constant probability. Because of our rejection sampling, this really requires obtaining $(b+d)\cdot 2^{O(d)}$ samples from $\mathcal{D}$. The final result is that we have a $b$-robust certificate of size only $O(b+d)$ for $x$. This required us to draw only a polynomially larger sample than we would have if we directly applied Theorem 5.2.

### v4

@@ 3 Worst Case Bounds on Certificate Size
Let $\{(x_{1},y_{1}),\dots,(x_{s_{b}},y_{s_{b}})\}$ be the largest $b$-robust hollow star for $\mathcal{H}$, with corresponding weights $w_{1},\dots,w_{s_{b}}$. Let $i^{*}$ be such that $w_{i^{*}}=b+1$. We will set the training dataset $S=\{(x_{i},y_{i}):i\neq i^{*}\}$, test point to be $x_{i^{\star}}$ and test label to be $1-y_{i^{\star}}$, and argue that $S$ is a minimal $b$-robust certificate for $(x_{i^{\star}},1-y_{i^{\star}})$.
Let $T$ be the weighted sequence $\{(x_{1},y_{1},w_{1}),\dots,(x_{s_{b}},y_{s_{b}},w_{s_{b}})\}$. By definition of the $b$-robust hollow star, $T^{\prime}=T\setminus\{(x_{i^{*}},y_{i^{*}},w_{i^{*}})\}$ is $b$-robustly realizable by $\mathcal{H}$. But note that all the weights in $T^{\prime}$ are equal to $1$. This means that there exists $h\in\mathcal{H}$ such that $\sum_{i\in[s_{b}],i\neq i^{*}}\mathds{1}[h(x_{i})\neq y_{i}]\leq b$. Namely, $S$ is $b$-robustly realizable by $\mathcal{H}$.
Next, consider any $h\in\mathcal{H}$ that satisfies $\sum_{i\in[s_{b}],i\neq i^{*}}\mathds{1}[h(x_{i})\neq y_{i}]\leq b$. Then, it must be the case that $h(x_{i^{*}})=1-y_{i^{*}}$; otherwise, the weighted sequence
$T=\{(x_{1},y_{1},w_{1}),\dots,(x_{i^{*}},y_{i^{*}},w_{i^{*}}),\dots,(x_{s_{b}},y_{s_{b}},w_{s_{b}})\}$
would be $b$-robustly realizable by $\mathcal{H}$, which would contradict that $\{(x_{1},y_{1}),\dots,(x_{s_{b}},y_{s_{b}})\}$ is a $b$-robust hollow star. Thus, $S$ is a $b$-robust certificate for $(x_{i^{*}},1-y_{i^{*}})$.
@@
Let us revisit the examples of halfspaces and singletons considered earlier for the setting with no corruptions. For $d$-dimensional halfspaces passing through the origin (Example 3.1), our certificate based on Carathéodory’s theorem had size $d$. Indeed, this class has hollow star number equal to $d+1$, which means that this certificate size is optimal in general. On the other hand, the class of singletons on a domain of size $n$ has hollow star number equal to $n$. This validates why we were unable to obtain a certificate of size smaller than n1n-1 in Example 3.2.
Thus, we have shown that in general, the $b$-robust hollow star number optimally characterizes minimum reliable certificate size. While Claim 3.6 relates the $b$-robust hollow star to the hollow star number (which is known for many natural classes, e.g., see the examples in Section 2.1 in BHMZ (20)), it merely gives a lower bound. It would be interesting to see if an upper bound can also be obtained, and more generally to chart out the $b$-robust hollow star for natural classes.
For example, we can show that the $b$-robust hollow star number for singletons on a domain of size $n$ is exactly (b+1)(n1)+1(b+1)(n-1)+1 (Section A.2). In addition, for $d$-dimensional halfspaces, one can use the same reasoning as in Example 3.1, together with the Tolerance Carathéodory theorem (Thereom 4.1 in MO (11), Theorem 2 in Tuz (89)) to obtain certificates of size $(d+b)^{O(b)}$ for test points in the $b$-robust agreement of the training data. From Theorem 3.7, we then know that the $b$-robust hollow star number of halfspaces is at most $(d+b)^{O(b)}$ (the lower bound from Claim 3.6 is only $\Omega(bd)$.) One way to close the gap for halfspaces would be to show a lower bound for the Tolerance Carathéodory theorem. The simplest phrasing of this for $b=1$ amounts to the following purely convex-analytic question:
@@
Such a result exists for the Tolerance Helly Theorem (MO, 11, Theorem 3.2).
Update:
We were recently made aware that the answer to the question above is affirmative, as established by Bárány and Perles: see Theorems 3 and 4 in BP (90). In particular, for fixed $b$ and as $d\to\infty$, the $b$-robust hollow star number of $d$-dimensional halfspaces is lower-bounded by $\frac{1}{e^{b}(b+1)!}d^{b+1}$.
@@ 4 Worst-case Distributional Bounds on Certificate Size
A first observation is that if the distribution $\mathcal{D}$ is discrete, and the test point $x$ belongs to the support of $\mathcal{D}$, we will eventually observe more than $2b+1$ copies of it in $S$. With a corruption budget of $b$, an adversary can potentially corrupt the labels on $b$ of these copies; however, $b+1$ copies with the true label $h^{\star}(x)$ still remain in $S$. We can use these copies, all of which have the true label, to certify that $x$ should be labeled as $h^{\star}(x)$, since any hypothesis that labels $x$ differently makes strictly more than $b$ mistakes on these copies. Furthermore, this certificate size is optimal—any certificate of smaller size is $b$-robustly realizable by every hypothesis in the class.
@@
Let $m^{\prime}$ be such that $\operatorname*{\mathbb{P}}_{S\sim\mathcal{D}^{m^{\prime}}}[(x,h^{\star}(x))\text{ in agreement region of }S]=\delta^{\prime}$ for some $\delta^{\prime}>0$; such an $m^{\prime}$ exists by assumption that $(x,h^{\star}(x))$ belongs to the agreement region of a drawn sample $S$ eventually. Consider drawing a sample $S$ of size $(2b+1)nm^{\prime}/\delta^{\prime}$, and let $S_{1},S_{2},\dots,S_{(2b+1)n/\delta^{\prime}}$ denote independent partitions of $S$ into chunks of size $m^{\prime}$. The expected number of chunks that satisfy that $(x,h^{\star}(x))$ is in the agreement region of the chunk is $(2b+1)n$. By a Chernoff bound, the probability that at least $2b+1$ chunks have $(x,h^{\star}(x))$ in their agreement region is at least $1-e^{-\Omega(bn)}$. Now, with a corruption budget of size $b$, an adversary can insert a corruption in at most $b$ chunks. Even so, the remaining $\geq b+1$ chunks continue to have $(x,h^{\star}(x))$ in their agreement region. Without loss of generality, suppose $S_{1},\dots,S_{b+1}$ are all uncorrupted. Then, from Theorem 3.7 (with $b=0$), we know that each of $S_{1},\dots,S_{b+1}$ will contain a certificate for $(x,h^{\star}(x))$ of size at most $s_{0}-1$; let these certificates be $C_{1},\dots,C_{b+1}$. The final $b$-robust certificate is simply a concatenation of $C_{1},\dots,C_{b+1}$.
@@
and this $h$ must label $x_{i}$ as $y_{i}$ (by non-realizability of the hollow star). Then, because the certificate only consists of at most $b$ copies of $(x_{i},y_{i})$, it is $b$-robustly realizable by $h$. This contradicts it being a $b$-robustly realizable certificate for $(x_{s_{0}},-y_{s_{0}})$, since $h$ labels $x_{s_{0}}$ as $y_{s_{0}}$.
@@ 5.1 Upper Bound on the Sample Complexity
We observe that the proof of Theorem 5.2 can be adapted to ensure that $S$ robustly certifies multiple test points simultaneously. Concretely, for a class $\mathcal{H}$ having VC dimension $d$, target hypothesis $h^{\star}\in\mathcal{H}$, and marginal distribution $\mathcal{D}$, let $\mathcal{X}_{\varepsilon}=\{x\in\mathcal{X}:\varepsilon_{x}(\mathcal{D},h^{\star})\geq\varepsilon\}$. Observe that any $h$ that mislabels at least one point—say $x_{1}$—in $\mathcal{X}_{\varepsilon}$ satisfies that $h\in\mathcal{H}_{x_{1}}\implies\operatorname*{\mathbb{P}}_{z\sim\mathcal{D}}[h(z)\neq h^{\star}(z)]\geq\varepsilon$ by definition of $\varepsilon_{x_{1}}$. From here, the analysis in the proof of Theorem 5.2 gives us that for a sample $S$ of size $O\left(\frac{b+d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right)$, with probability $1-\delta$, every point in $\mathcal{X}_{\varepsilon}$ belongs to the $b$-robust agreement region of SS. [^3] One can also analyze the probability mass of $\mathcal{X}_{\epsilon}$, as it is directly related to the disagreement coefficient $\theta$ introduced by Han (07). Specifically, it holds that $\operatorname*{\mathbb{P}}[x\in\mathcal{X}_{\varepsilon}]=1-\operatorname*{\mathbb{P}}[x\in\mathcal{X}\setminus\mathcal{X}_{\varepsilon}]\geq 1-\theta\varepsilon$.
@@ 5.2 Tightness of the Upper Bound
To show the tightness of the other two terms, we will work with the following class $\mathcal{H}$. For any given $d\geq 1$, the domain $\mathcal{X}$ of the class will be $\{x_{0},\dots,x_{k}\}$, where we can pick $k$ to be any integer that satisfies $1+\sqrt{\frac{100}{d}}<\frac{1}{2}\log\left(\frac{k}{d}\right)$. [^4] This is for the purpose of instantiating a coupon collector bound, e.g., as in Remark 20, BHMZ (20). For example, $k=(100d)^{20}$ is a valid choice. The hypothesis class $\mathcal{H}$ comprises of $\binom{k}{d}+1$ hypotheses. For every $S\subset\{x_{1},\dots,x_{k}\},|S|=d$, there is a hypothesis $h_{S}\in\mathcal{H}$ that labels exactly the set $S$ as 1, and the rest of the domain as 1-1. Additionally, there is a special hypothesis $h^{\star}\in\mathcal{H}$ which satisfies that $h^{\star}(x_{0})=1,h^{\star}(x_{i})=-1,\forall i\in[k]$. That is, $h^{\star}$ only labels $x_{0}$ as 1. Moreover, $h^{\star}$ is the only hypothesis in $\mathcal{H}$ that labels $x_{0}$ as $1$; all the other $\binom{k}{d}$ hypotheses label $x_{0}$ as 1-1. It can be readily verified that the VC dimension of $\mathcal{H}$ is $d$. In what follows, we will set $b=0$ for convenience.
We will set the target hypothesis to be $h^{\star}$, the data distribution $\mathcal{D}$ to be uniform on $\{x_{1},\dots,x_{k}\}$, and the test point $x$ to be $x_{0}$. Crucially, $\mathcal{D}$ has zero mass on $x_{0}$.  [^5] We can also allow a mass of $o(\varepsilon_{x}/(d\log(1/\varepsilon_{x})))$ on $x_{0}$—this ensures $x_{0}$ is not seen in the sample with good chance. Recall that $h^{\star}(x_{0})=1$ and $h^{\star}(x_{i})=-1,\forall i\in[k]$. Since every other hypothesis in $\mathcal{H}$ labels $x_{0}$ as 1-1, and also labels some $d$ points in $\{x_{1},\dots,x_{k}\}$ as 1, we have that $\varepsilon_{x}=\frac{d}{k}$. Now suppose that we obtain a sample of size $m=\frac{d}{2\varepsilon_{x}}\log\left(\frac{1}{\varepsilon_{x}}\right)=\frac{k}{2}\log\left(\frac{k}{d}\right)$. Then, by a coupon collector argument (see Lemma 19 in BHMZ (20)), we have that with probability at least $\frac{1}{100}>\delta$, the number of distinct elements from $\{x_{1},\dots,x_{k}\}$ that we see in the sample are at most kdk-d. But this means that there is some subset $S\subset\{x_{1},\dots,x_{k}\}$ of $d$ elements whose labels we have not seen. Thus, $h_{S}$ is consistent with the labeled data seen so far, rendering $(x,h^{\star}(x))$ to not be in the agreement region of the sample.
@@ 6 Shorter Certificates by Reweighting
Consider the case when $\mathcal{H}$ is the class of halfspaces in $\mathbb{R}^{d}$ (not necessarily passing through the origin), and $\mathcal{D}$ is uniform on the unit ball $\{x:\|x\|_{2}\leq 1\}$. Suppose that the target halfspace $h^{\star}$ labels every point in the ball positively (i.e., it does not cut through the ball), and that the test point $x$ is at $(1/2,0,0,\dots,0$). We can verify that $\varepsilon_{x}\leq 2^{-\Omega(d)}$ (as realized by the halfspace $x_{1}<1/2$; see Theorem 2.7 in BHK (20) for the volume of the ball excluded by this halfspace), and so, the upper bound in Theorem 5.2 would suggest that we draw $b\cdot 2^{\Omega(d)}$ samples to ensure that $x$ lies in the $b$-robust agreement region of the samples with constant probability. We might extract a shorter certificate from this sample using the worst-case bound from Theorem 3.7, but even this scales as the $b$-robust hollow star number of $\mathcal{H}$, for which the best upper bound is $(d+b)^{O(b)}$. Consider the following alternate recipe. Suppose we discard any samples from $\mathcal{D}$ that are not contained in the ball of radius $1/2$ centered at $x$. This induces the uniform distribution $\mathcal{D}_{w}$ on the smaller ball, and in expectation, we obtain one sample from $\mathcal{D}_{w}$ for every $2^{d}$ samples from $\mathcal{D}$. Crucially, notice that $\varepsilon_{x}(\mathcal{D}_{w},h^{\star})=1/2$. By Theorem 5.2, $O(b+d)$ samples from $\mathcal{D}_{w}$ are sufficient to ensure that $x$ lies in the $b$-robust agreement region of the samples with constant probability. Because of our rejection sampling, this really requires obtaining $(b+d)\cdot 2^{O(d)}$ samples from $\mathcal{D}$. The final result is that we have a $b$-robust certificate of size only $O(b+d)$ for $x$. This required us to draw only a polynomially larger sample than we would have if we directly applied Theorem 5.2.